OpenStudy (anonymous):

1. Find the equation for a hyperbola centered at (0, 0), with foci at (0, -) and (0, ) and vertices at (0, -8) and (0, 8). 2. Find the equation for a parabola with its focus at (-7, 0) and a directrix of x = 7. 3. Find the equation of a circle with its center at (2, 7) and a radius of 5.

OpenStudy (anonymous):

\[y^2-x^2=1\]

OpenStudy (anonymous):

1) I think you missng a

OpenStudy (anonymous):

3) \[(x-2)^2+(y-7)^2=5^2\]